# Risk-Budgeted Control Framework for Balanced Performance and Safety in Autonomous Vehicles

### v1

@@ IV-A Bad Step detection
$b_{k}=\mathbf{1}\{\nu_{k}>\bar{\nu}\}\ \lor\ \mathbf{1}\{r_{\min,k}<\delta\},\quad b_{k}\in\{0,1\},$ (19)
@@
IV-B Sliding Window Risk-Budgeted Monitor
@@
$h_{k}\geq\mu^{W}h_{k}+c\sum_{j=0}^{W-1}\mu^{W-1-j}v_{k-j}.$ (20)
Here, $v_{k-j}$ is a variable representing the worst-case lower bound of the residual $r_{k-j}$ used in the proof. Thus, $v_{k-j}$ takes a value from the set $\{\delta,-\overline{\nu}\}$. This is because a good step is guaranteed to have a residual of at least $\delta$ (in section IV-A), while a bad step has its residual lower-bounded by $\overline{\nu}$. To guarantee safety ($h_{k}\geq 0$), we must find a lower bound for this expression. The core challenge lies in finding the worst-case arrangement of the $M$ bad steps in the summation. The next lemma establishes that the worst case occurs at the end of the step. Consequently, we can apply the rearrangement lemma to determine the parameters in this worst-case configuration.
@@
Let the weights $w_{j}:=\mu^{W-1-j}$ for $j=0,\dots,W-1$. Since $\mu\in(0,1)$, these weights are monotonically increasing: $0<w_{0}<w_{1}<\dots<w_{W-1}$. Let $V$ be the set of sequences $\{v_{j}\}_{j=0}^{W-1}$ with the worst case $M$ entries equal to $\overline{\nu}$ and WMW-M entries equal to $\delta$. The weighted sum $S(v):=\sum_{j=0}^{W-1}w_{j}v_{j}$ is minimized over $v\in V$ by assigning the $M$ negative values $(-\overline{\nu})$ to the largest weights and the WMW-M positive values $(+\delta)$ to the smallest weights.
@@
Consider any pair of indices $i<j$, which implies $w_{i}\leq w_{j}$. Let’s assume an arrangement where $v_{i}=\delta$ and $v_{j}=-\overline{\nu}$. The contribution from this pair to the sum is $w_{i}\delta+w_{j}(-\overline{\nu})$. By swapping the values: $v^{\prime}_{i}=-\overline{\nu}$ and $v^{\prime}_{j}=\delta$. The new contribution is $w_{i}(-\overline{\nu})+w_{j}\delta$. The difference between the original sum and swapped sum is:
@@
Applying the Rearrangement Lemma to the summation in Equation (20), we can establish the definitive worst-case lower bound for $h_{k}$. The minimum is achieved when the $M$ bad steps are assigned to the terms with the largest weights (corresponding to indices $j=W-M,\dots,W-1$) and the WMW-M good steps are assigned to the terms with the smallest weights (indices $j=0,\dots,W-M-1$).
@@
$h_{k}\geq\mu^{W}h_{k}+c\left[\delta\sum_{l=M}^{W-1}\mu^{l}-\overline{\nu}\sum_{l=0}^{M-1}\mu^{l}\right].$ (21)
@@
Assume the one-step comparison inequality (9) holds with $0<\mu<1$, $c>0$. In any sliding window of length $W$, there are at most $M$ bad steps (with $r\geq-\overline{\nu}$) and WMW\!-\!M good steps (with $r\geq\delta$). Moreover, each window starts from a safe boundary state $h_{k}\geq 0$. Then the terminal state is safe ($h_{k}\geq 0$) whenever
@@
To ensure safety, we require that the lower bound of $h_{k}$ remains non-negative. Starting from the worst-case inequality established in (21) and conservatively dropping the $\mu^{W}h_{k}\geq 0$ term, the condition becomes:
@@
Rather than forbidding any instantaneous violation, the certificate caps how many bad steps and how severe they can be via $(W,M,\delta,\overline{\nu})$, thereby guaranteeing $h_{k}\geq 0$ at every window boundary. Once the $W,M,\delta$ has been designed based on the system requirement, we can get $\overline{\nu}$ from the equation below:
@@
V Risk-Switch for Probabilistic Safety
Building on the risk monitor in Section IV, we formalize a sliding-window supervisor that certifies safety at the window scale and coordinates switching from the performance-seeking R-CBF filter to a conservative RC-CBF safety filter when risk accumulates. The cumulative number of bad steps $m_{k}$ in the most recent $W$-step window is updated by
@@ V Risk-Switch for Probabilistic Safety
V-A Feasibility-Triggered CVaR-CBF (FT-C-CBF)
Let $\mathcal{A}$ denote the feasible set of the R-CBF filter at time $k$, and let $a_{k}=\mathbf{1}\{(u_{k},\nu_{k})\in\mathcal{F}\}$ be its feasibility flag. Under FT, feasibility is the primary guard, while the sliding-window counter enforces a budget on accumulated risk:
If the R-CBF problem is infeasible at time $k$ ($a_{k}=0$), the monitor updates $b_{k}$ and $m_{k}$ via (26). When the cumulative count reaches the budget $M$ (i.e., $m_{k}\geq M$), the monitor switches from the R-CBF filter to the CVaR-CBF controller.
$\eta_{k}=\begin{cases}1,&\text{if }(a_{k}=0)\ \land\ (m_{k}\geq M),\\ 0,&\text{if }(a_{k}=1)\ \land\ (m_{k}<M),\end{cases}$
with $\eta_{k}\in\{0,1\}$ the controller mode (0: R-CBF, $1$: CVaR-CBF). The controller $\eta_{k}$ will be implemented at time $k$.
V-B Quality-Triggered CVaR-CBF (QT-C-CBF)
In the QT mechanism, every step is vetted after solving the R-CBF filter, irrespective of feasibility, the risk monitor computes $b_{k}$ and updates the sliding-window count $m_{k}$. Switching depends solely on whether the window-level risk budget is exceeded.
@@ V-B Quality-Triggered CVaR-CBF (QT-C-CBF)
Thus, even when the R-CBF problem is feasible, a step flagged as bad by the monitor contributes to $m_{k}$; once the cumulative count reaches the budget $M$, the supervisor switches from the R-CBF filter to the CVaR controller.
@@
VI Simulations
VI-A Experimental Setup
VI-A1 Environment
We evaluate a kinematic vehicle under an MPC controller, following the experimental setup of [1]. The reference trajectory is converted from a GPS path collected from the experimental autonomous vehicle reported in [1] and validated.
VI-A2 Safety Framework Configuration
The minimum safety distance set to $D_{s}$=3m, and $\kappa=1$, and the sliding window risk monitor is configured with a window size of W=5, a maximum of M=1 bad step, and a good step safety margin $\delta=1$m. Based on these parameters, the risk cap is calculated using (25) to be $\overline{\nu}$= 3.8.
The Conditional Value-at-Risk (CVaR) uses a confidence level of $\epsilon$=95% and a loss threshold of $\gamma$=0, consistent with the definition in (17). To simulate realistic localization uncertainty, the vehicle’s position is modeled as a Gaussian distribution. At each timestep, Q=10 samples are drawn from the distribution, which is centered on the vehicle’s measured position and has a standard deviation $\sigma_{v}=0.1$, to represent the set of possible true locations.
VI-A3 Obstacle Uncertainty Modeling
To account for obstacle movement and perception errors, we model uncertainty in the obstacle’s state. At each timestep $k$, upon receiving an obstacle’s measured position $\bm{x}_{m,k}^{o}$, we generate P=10 samples to represent its possible locations $\bm{x}_{k}^{o,p}$. These samples are uniformly distributed within a square bounding box of half width $\sigma_{o}\in\{1,2,3,4,5\}$m centered on the measured position. This setup simulates a perception detection that provides a bounded box estimate of the obstacle’s location, with the chosen parameters representing an uncertainty scenario.
@@ VI-A3 Obstacle Uncertainty Modeling
$\Delta\bm{x}_{k}^{o,p}\sim\mathcal{U}\left(\left[-{\sigma_{o}},{\sigma_{o}}\right]^{2}\right).$
VI-A4 Performance Metrics
To evaluate the controller’s performance, we define five metrics to aggregate performance over all test cases. Some of our aggregate metrics are adapted from the methodology presented in [14].
To assess overall performance across all $N_{t}$ test cases, we calculate the following metrics. Note that for metrics other than success rates, the averages are computed only over the set of successful runs, $N_{s}$.
  • Success Rate (SR): A run is successful if the minimum vehicle pedestrian distance along the trajectory always exceeds the required clearance $d_{\min}=2.8\,\mathrm{m}$ (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$). The success rate is the fraction of successful runs: $\mathrm{SR}=N_{s}/N_{t}$.
  • Minimum Distance to Pedestrian (MDP): For each successful run, let $d_{m,i}$ be the minimum distance to the nearest pedestrian while avoiding; we report the mean over successful runs, $MDP=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}d_{m,i}$. An ideal clearance of at least $5.8\,\mathrm{m}$ is required (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$ + safety buffer $3.0\,\mathrm{m}$). Values $\geq 5.8\,\mathrm{m}$ satisfy the safety margin; smaller values indicate the controller is performance-based to track the reference path as close as possible.
  • Infeasibility Rate (IR): The proportion of optimization steps that failed to yield a feasible solution, calculated as $IR_{i}=n_{i}/n_{t}$, where $n_{i}$ is the number of infeasible steps and $n_{t}$ is the total number of optimization steps. The mean infeasibility solution rate across all successful runs, $IR=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}IR_{i}$.
  • Calculation Time (CT): The mean calculation time for each step in all successful runs, $CT=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}T_{i}$.
  • Cross-Track Error (CTE): To evaluate path-following performance, we calculate the deviation from the desired path. We record the Average CTE ($E_{i}$), calculated as the mean of all CTE measurements $e_{k}$ during the run. The mean of the average CTE values from all successful runs, $CTE=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}E_{i}$.
VI-B Comparative Analysis
@@
    Relaxed CBF (R-CBF): We enforce the CBF constraint via the relaxed formulation in (10) over the entire trajectory, using the raw, noisy pedestrian measurements as inputs to the filter (i.e., no stochastic pedestrian and vehicle states in the controller).
@@ VI-C Delta Comparison
VII Conclusion
We presented a control framework with a risk monitor to switch the safety filter between the performance CBF and conservative CVAR in autonomous driving scenarios. Monte–Carlo evaluations with localization and obstacle state uncertainty show that using the QT option in the control framework has a high success rate ($\geq\!95\%$) and maintains low CTE showing the trajectory following performance while avoiding the pedestrians. When safety quality is high, the risk monitor allow the controller in its performance mode; as the margin degrades, the monitor reliably triggers the CVaR controller to preserve safety throughout the trajectory. The proposed framework achieves a performance safety balanced with risk-budgeted guarantees and potential computation compatible with on–board control loops. As next steps, we will deploy the framework under the proposed risk-budgeted monitor on our autonomous vehicle platform for hardware–in–the–loop and on–road validation.

### v2

@@ III-A Bad Step detection
$b_{k}=\mathbf{1}\{r_{k}<\delta\},\quad b_{k}\in\{0,1\},$ (15)
@@
III-B Sliding Window Risk-Budgeted Monitor
@@
$h_{k+1}\geq\mu^{W}h_{k+1-W}+c\sum_{j=0}^{W-1}\mu^{W-1-j}\tilde{r}_{k-j}.$ (16)
Here, $\tilde{r}_{k-j}$ is a variable representing the worst-case lower bound of the residual $r_{k-j}$ used in the proof. Thus, $\tilde{r}_{k-j}$ takes a value from the set $\{\delta,-\overline{\nu}\}$. This is because a good step is guaranteed to have a residual of at least $\delta$ (in section III-A), while a bad step has its residual lower-bounded by $\overline{\nu}$. To guarantee safety ($h_{k+1}\geq 0$), we must find a lower bound for this expression. The core challenge lies in finding the worst-case arrangement of the $M$ bad steps in the summation. The next lemma establishes that the worst case occurs at the end of the step. Consequently, we can apply the rearrangement lemma to determine the parameters in this worst-case configuration.
@@
Fix any $k$ and define the window-indexed sequence $\{\tilde{r}_{j}\}_{j=0}^{W-1}$ by $\tilde{r}_{j}:=\tilde{r}_{k-j}$ for $j=0,\dots,W-1$. Let the weights $w_{j}:=\mu^{W-1-j}$ for $j=0,\dots,W-1$. Since $\mu\in(0,1)$, these weights are monotonically increasing: $0<w_{0}<w_{1}<\dots<w_{W-1}$. Let $V$ be the set of sequences $\{\tilde{r}_{j}\}_{j=0}^{W-1}$ with the worst case $M$ entries equal to $\overline{\nu}$ and WMW-M entries equal to $\delta$. The weighted sum $S(\tilde{r}):=\sum_{j=0}^{W-1}w_{j}\tilde{r}_{j}$ is minimized over $\tilde{r}\in V$ by assigning the $M$ negative values $(-\overline{\nu})$ to the largest weights and the WMW-M positive values $(+\delta)$ to the smallest weights.
@@
Consider any pair of indices $i<j$, which implies $w_{i}\leq w_{j}$. Let’s assume an arrangement where $\tilde{r}_{i}=\delta$ and $\tilde{r}_{j}=-\overline{\nu}$. The contribution from this pair to the sum is $w_{i}\delta+w_{j}(-\overline{\nu})$. By swapping the values: $\tilde{r}^{\prime}_{i}=-\overline{\nu}$ and $\tilde{r}^{\prime}_{j}=\delta$. The new contribution is $w_{i}(-\overline{\nu})+w_{j}\delta$. The difference between the original sum and swapped sum is:
@@
Applying the Rearrangement Lemma to the summation in Equation (16), we can establish the definitive worst-case lower bound for $h_{k+1}$. The minimum is achieved when the $M$ bad steps are assigned to the terms with the largest weights (corresponding to indices $j=W-M,\dots,W-1$) and the WMW-M good steps are assigned to the terms with the smallest weights (indices $j=0,\dots,W-M-1$).
@@
$h_{k+1}\geq\mu^{W}h_{k+1-W}+c\left[\delta\sum_{l=M}^{W-1}\mu^{l}-\overline{\nu}\sum_{l=0}^{M-1}\mu^{l}\right].$ (17)
@@
Assume the one-step comparison inequality (5) holds with $0<\mu<1$ and $c>0$. Suppose the safety filter enforces the uniform lower bound $r_{t}\geq-\overline{\nu}$ at every step $t$. Fix any $k$ and consider the window of length $W$ ending at $k$, i.e., steps $t=k+1-W,\dots,k$. Assume that at most $M$ steps in this window are bad (i.e., satisfy $r_{t}<\delta$), and that the window starts from a safe state $h_{k+1-W}\geq 0$. Then the terminal state at the end of the window is safe, i.e., $h_{k+1}\geq 0$, whenever
@@
To ensure safety, we require that the lower bound of $h_{k+1}$ remains non-negative. Starting from the worst-case inequality established in (17) and conservatively dropping the $\mu^{W}h_{k+1-W}\geq 0$ term, the condition becomes:
@@
Rather than forbidding any instantaneous violation, the certificate caps how many bad steps and how severe they can be via $(W,M,\delta,\overline{\nu})$, thereby guaranteeing $h_{k+1}\geq 0$ at every window boundary. Once the $W,M,\delta$ has been designed based on the system requirement, we can get $\overline{\nu}$ from the equation below:
@@
The sliding-window certificate bounds the accumulation of bad steps and the resulting decay of the barrier value under the assumption that the system starts within the safe set, i.e., $h(x_{0})\geq 0$, and that feasible control actions exist for the considered constraints. Accordingly, the proposed monitoring mechanism is preventive. It regulates how often safety margin violations may occur so as to avoid leaving the safe set. If the system state enters a region with $h(x)<0$ where no admissible input can drive it back to $h(x)\geq 0$ (due to actuation limits or environment geometry), then switching controllers cannot restore invariance and recovery is not guaranteed.
IV Risk-Switch for Probabilistic Safety
Building on the risk monitor in Section III, we formalize a sliding-window supervisor that certifies safety at the window scale and coordinates switching from the performance-seeking R-CBF filter to a conservative CVaR-CBF safety filter when risk accumulates. The cumulative number of bad steps $m_{k}$ in the most recent $W$-step window is updated by
@@ IV Risk-Switch for Probabilistic Safety
IV-A Feasibility-Triggered CVaR-CBF (FT-C-CBF)
To handle non-optimal outcomes in a unified way, we use the solver’s feasibility status as a trigger. At 20–50 Hz, the solver often returns the same “infeasible” status for both true infeasibility and non-optimal termination, so we do not attempt to disambiguate online; instead, we treat it as a trigger and classify the step via the residual test $r_{t}<\delta$.
Let $\mathcal{F}_{k}$ denote the feasible set of the R-CBF filter at time $k$, and let $a_{k}=\mathbf{1}\{(u_{k},\nu_{k})\in\mathcal{F}\}$ be its feasibility flag. Under FT, feasibility serves as the primary guard, while the sliding-window counter enforces a budget on accumulated risk. If the solver under the R-CBF constraints reports infeasible at time $k$ ($a_{k}=0$), the monitor updates $b_{k}$ and $m_{k}$ via (22). When the solver reports infeasible at time $k$ ($a_{k}=0$), the monitor updates $b_{k}$ and $m_{k}$ via (22). If an infeasible report occurs when the cumulative count has reached the budget (i.e., $a_{k}=0$ and $m_{k}\geq M$), the monitor switches from the R-CBF filter to the CVaR-CBF controller:
$\eta_{k}=\begin{cases}1,&\text{if }(a_{k}=0)\ \land\ (m_{k}\geq M),\\ 0,&\text{otherwise},\end{cases}$
where $\eta_{k}\in\{0,1\}$ is the controller mode (0: R-CBF, $1$: CVaR-CBF). The mode $\eta_{k}$ is implemented at time $k$.
IV-B Quality-Triggered CVaR-CBF (QT-C-CBF)
The quality trigger is motivated by the observation that optimality alone does not guarantee sufficient future safety margin. Accordingly, at every step we check whether the residual satisfies the safety-margin requirement, and mark the step as bad whenever $r_{k}<\delta$. Under QT, each step is vetted after solving the R-CBF constraints: the risk monitor computes $b_{k}$ and updates the sliding-window count $m_{k}$. In contrast to the feasibility-triggered scheme, switching depends only on whether the window-level risk budget is exceeded:
@@ IV-B Quality-Triggered CVaR-CBF (QT-C-CBF)
Thus, even when the solver returns an optimal solution, a step flagged as bad by the monitor contributes to $m_{k}$; once the cumulative count reaches the budget $M$, the supervisor switches from the R-CBF filter to the CVaR controller.
@@
V Simulations
V-A Experimental Setup
V-A1 Environment
We evaluate a kinematic vehicle controlled by MPC following the experimental setup of [10]. The reference trajectory is obtained from a GPS path recorded by the experimental AV in [10].
V-A2 Safety Framework Configuration
The minimum safety distance is $D_{s}=3$ m and $\kappa=1$. The sliding-window monitor uses $W=5$, allowing at most $M=1$ bad step with a safety margin $\delta=1$ m. Based on these parameters, the risk cap computed from (21) is $\overline{\nu}=3.8$.
For the CVaR constraint, the confidence level is $\epsilon=95\%$ with loss threshold $\gamma=0$, consistent with (13). Vehicle localization uncertainty is modeled as a Gaussian distribution with standard deviation $\sigma_{v}=0.1$. At each timestep, $Q=10$ samples are drawn around the measured position to represent possible true vehicle states.
V-A3 Obstacle Uncertainty Modeling
To account for obstacle motion and perception uncertainty, the obstacle state is sampled at each timestep. Given the measured position $\bm{x}_{m,k}^{o}$, we generate $P=10$ samples $\bm{x}_{k}^{o,p}$ uniformly within a square bounding box centered at $\bm{x}_{m,k}^{o}$ with half width $\sigma_{o}\in\{1,2,3,4,5\}$ m. This models a perception system that reports a bounded box estimate of the obstacle location.
@@ V-A3 Obstacle Uncertainty Modeling
$\Delta\bm{x}_{k}^{o,p}$ $\sim\mathcal{U}\!\left([-\sigma_{o},\sigma_{o}]^{2}\right).$
V-A4 Performance Metrics
We evaluate performance using five aggregate metrics, adapted in part from [16]. Let $N_{t}$ denote the total number of test cases and $N_{s}$ the number of successful runs. Except for success rate, all metrics are averaged over successful runs only.
  • Success Rate (SR): A run is successful if the minimum vehicle–pedestrian distance exceeds $d_{\min}=2.8$ m throughout the trajectory. The success rate is $\mathrm{SR}=\frac{N_{s}}{N_{t}}.$
  • Minimum Distance to Pedestrian (MDP): Let $d_{m,i}$ denote the minimum distance to the nearest pedestrian in successful run $i$. We report $\mathrm{MDP}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}d_{m,i}.$
    A clearance of at least $5.8$ m satisfies the desired safety margin.
  • Infeasibility Rate (IR): For successful run $i$, let $IR_{i}=n_{i}/n_{t}$, where $n_{i}$ is the number of infeasible steps and $n_{t}$ is the total number of optimization steps. We report $\mathrm{IR}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}IR_{i}.$
  • Calculation Time (CT): The mean per-step computation time over successful runs: $\mathrm{CT}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}T_{i}.$
  • Cross-Track Error (CTE): Let $E_{i}$ denote the average cross-track error in successful run $i$. We report $\mathrm{CTE}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}E_{i}.$
V-B Comparative Analysis
@@
    Relaxed CBF (R-CBF, baseline): We apply the relaxed CBF filter (6) in a plug-in manner using the raw noisy pedestrian measurements as the state input. This baseline represents a common naïve deployment of deterministic CBF filters under sensing noise, where estimation uncertainty is not modeled in the constraint. The purpose is to quantify the robustness gap caused by ignoring uncertainty.
@@ V-C Delta Comparison
VI Conclusion
We presented a hybrid control framework with a risk monitor to switch the safety filter between the performance CBF and conservative CVAR in autonomous driving scenarios. Monte–Carlo evaluations with localization and obstacle state uncertainty show that using the QT option in the control framework has a high success rate ($\geq\!95\%$) and maintains low CTE showing the trajectory following performance while avoiding the pedestrians. When safety quality is high, the risk monitor allow the controller in its performance mode; as the margin degrades, the monitor reliably triggers the CVaR controller to preserve safety throughout the trajectory. The proposed framework achieves a performance safety balanced with risk-budgeted guarantees and potential computation compatible with on–board control loops. As next steps, we will deploy the framework under the proposed risk-budgeted monitor on our autonomous vehicle platform for hardware–in–the–loop and on–road validation.